OpenStudy (anonymous):

Write the equation of the function g(x) if g(x) = f(x – 3) +1 and f(x) = x^3 + 1

OpenStudy (campbell_st):

in f(x) just replace x with x - 3

OpenStudy (anonymous):

i got g(x)=((x^3+1)-3)+1 when i did that

OpenStudy (anonymous):

but i get stuck on what to do after

OpenStudy (campbell_st):

well f(x -3) = (x -3)^3+1 then g(x) = (x -2)^3 + 1 + 1 then simplify

OpenStudy (campbell_st):

oops should read g(x) = (x -3)^3 + 1 + 1 then simplify

OpenStudy (anonymous):

((x-3)^3+1)+1 Distribute (x-3)^3 ! ((x^3-9 x^2+27 x-27)+1)+1

OpenStudy (anonymous):

g(x) = x^3-9 x^2+27 x-25